OpenStudy (daenio):

Scooby-Doo has become very successful at fighting crime. He and Shaggy can successfully solve a case 80% of the time. a) What is the probability that Scooby and Shaggy will not solve the first four cases before being successful? [2] b) What is the probability that Scooby’s first three cases will be unsuccessful? [2]

OpenStudy (daenio):

@blockcolder @satellite73

OpenStudy (lgbasallote):

since 20% is the fail rate and in fraction form it is 1/5 i'm thinking \[\frac{1}{5} \times \frac{1}{5} \times \frac{1}{5} \times \frac{1}{5}\] feel free to correct me though @blockcolder @satellite73

OpenStudy (nottim):

I don't understand Chloro. what happened?

OpenStudy (anonymous):

It's geometric type!

OpenStudy (daenio):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

have you ever seen or worked with the geometric probability distribution before?

OpenStudy (daenio):

Not really. I've worked more on hypergeometric distributions than geometric distributions.

jimthompson5910 (jim_thompson5910):

Well to answer this question, you would need to use the geometric distribution as the hypergeometric distribution is a lot like the binomial distribution (but no replacements are made)

OpenStudy (daenio):

Is the formula that I'm supposed to use in this document?

OpenStudy (daenio):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

alright one sec while I read the attachment

jimthompson5910 (jim_thompson5910):

It looks like you are using the formula you highlighted in bold on the left because that matches with the formula I'm looking at (the only difference is that q is 1 - p, but they are the same)

jimthompson5910 (jim_thompson5910):

Part a) Use the geometric probability distribution with p = 0.8 and k = 5 P(X = k) = (1-p)^(k-1)*p P(X = 5) = (1-0.8)^(5-1)*0.8 P(X = 5) = (0.2)^(4)*0.8 P(X = 5) = 0.0016*0.8 P(X = 5) = 0.00128 So the probability of the first success on the fifth trial (ie they solve their first case on the fifth try) is 0.00128

OpenStudy (daenio):

P(x)= q^x p = 0.20^4 x 0.80 = 0.00128

jimthompson5910 (jim_thompson5910):

You nailed it. Nice job

OpenStudy (daenio):

Hey, thank you! :)

jimthompson5910 (jim_thompson5910):

You're welcome. Want to take a stab at part b?

OpenStudy (daenio):

Sure!

OpenStudy (daenio):

P(x)= q^x p =0.20^3 x 0.80 =0.0064

jimthompson5910 (jim_thompson5910):

After looking at this part again in a bit more depth, I'm thinking that it's better modeled by the negative binomial probability distribution. Have you dealt with this distribution before?

jimthompson5910 (jim_thompson5910):

hmm nevermind that last thought. The negative binomial distribution is used to count the number of events and not the order of the events, so that's not too useful. lol sry, it's getting late over here and I'm falling asleep...

OpenStudy (daenio):

It's alright! Thanks for the help though!

OpenStudy (daenio):

How many cases would you expect Scooby and Shaggy to not solve before they are successful? [2] @satellite73

OpenStudy (daenio):

@jim_thompson5910

OpenStudy (daenio):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

In this part, you're finding the expected value of the geometric distribution, which is (1-p)/p where p is the probaility of success.

jimthompson5910 (jim_thompson5910):

probability*

OpenStudy (daenio):

1-0.80/ 0.80 = 0.20/0.80 = 0.25 @jim_thompson5910

jimthompson5910 (jim_thompson5910):

you got it, very nice work

OpenStudy (daenio):

Thanks. Hey, do you know how the equation begins?

jimthompson5910 (jim_thompson5910):

what do you mean?

OpenStudy (daenio):

Like for the formula, does it begin with "P(x)=" ?

jimthompson5910 (jim_thompson5910):

oh, it's just E(X) or E[X]

jimthompson5910 (jim_thompson5910):

to denote expected value

OpenStudy (daenio):

Oh ok. thanks!

jimthompson5910 (jim_thompson5910):

yw

OpenStudy (daenio):

Oh, but does that answer the question? How many cases would you expect Scooby and Shaggy to not solve before they are successful? [2] I would expect Scooby and Shaggy to not solve 0.25 cases before being successful?

jimthompson5910 (jim_thompson5910):

yes, that's exactly what the expected value finds

OpenStudy (daenio):

Oh ok.